OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2

OpenStudy (texaschic101):

x - y = 10 ---> x = y + 10 now sub y + 10 in for x in the second equation. 2x + y = 2 2(y + 10) + y = 2 Distribute the 2 through the parenthesis. Next, combine like terms. Get y on one side and everything else on the other side. Solve for y. Once you find y, you can sub that in to one of the original equations to solve for x. You can then check by subbing in your known variables and if both sides equal each other, then you have solved the equation correctly. Can you take it from here ?